### v1

# Reasoning via Video: The First Evaluation of Video Models’ Reasoning Abilities through Maze-Solving Tasks

Abstract
Video Models have achieved remarkable success in high-fidelity video generation with coherent motion dynamics. Analogous to the development from text generation to text-based reasoning in language modeling, the development of video models motivates us to ask: Can video models reason via video generation? Compared with the discrete text corpus, video grounds reasoning in explicit spatial layouts and temporal continuity, which serves as an ideal substrate for spatial reasoning. In this work, we explore the reasoning via video paradigm and introduce VR-Bench—a comprehensive benchmark designed to systematically evaluate video models’ reasoning capabilities. Grounded in maze-solving tasks that inherently require spatial planning and multi-step reasoning, VR-Bench contains 7,920 procedurally generated videos across five maze types and diverse visual styles. Our empirical analysis demonstrates that SFT can efficiently elicit the reasoning ability of video model. Video models exhibit stronger spatial perception during reasoning, outperforming leading VLMs and generalizing well across diverse scenarios, tasks, and levels of complexity. We further discover a test-time scaling effect, where diverse sampling during inference improves reasoning reliability by 10–20%. These findings highlight the unique potential and scalability of reasoning via video for spatial reasoning tasks.
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author.
Figure 1: Overview of VR-Bench. (A) Maze Types. VR-Bench comprises five maze types—Regular Maze, Irregular Maze, 3D Maze, Trapfield, and Sokoban—covering both 2D and 3D settings as well as diverse task structures, yielding a broad range of spatial reasoning scenarios. (B) Reasoning via Video Paradigm. VR-Bench adopts a chain-of-frame reasoning paradigm [wiedemer2025video], requiring models to produce frame-by-frame inferences that capture sequential visual reasoning. (C) Benchmark Performance. Leading VLMs and video models are evaluated on four core metrics across all maze types, revealing clear differences in spatial reasoning capability. (D) Additional Analysis. VR-Bench also supports evaluations on difficulty generalization, texture generalization, maze-type generalization, and test-time scaling, enabling a comprehensive assessment of model robustness and generalization.
1 Introduction
With the rapid development of diffusion-based and autoregressive-based generative architectures, video models have witnessed tremendous success in high-fidelity video generation. Previous works, such as Stable Video Diffusion [blattmann2023stable] and Imagen Video [ho2022imagen], showcase the capability of video models to generate physically realistic and temporally consistent videos conditioned on their input instructions. Recent studies further reveal that advanced video models are capable of performing a diverse range of visual tasks beyond generation itself, including perception, understanding, and even reasoning. These findings suggest that video models are evolving from pure generative models into general-purpose visual intelligence models. Analogous to the evolution of language models from text generation to text-based reasoning, the development of video models leads to a question: “Can video models reason via video generation?”
Crucially, the spatiotemporal nature of video modality offers a new perspective on reasoning. The traditional paradigm, which we term reasoning via text, uses language as the medium for expressing intermediate reasoning steps. Representative work, such as Chain-of-Thought prompting [wei2022chain, zhang2024aflow, yang2025multi, yuthought, wan2025deepresearch, yu2025recode, yao2023tree], achieves this by eliciting large language models (LLMs) to generate a coherent textual reasoning chain. Recently, this reasoning via text paradigm has been introduced to visual domains, including multimodal question answering and video understanding. However, even in these multimodal settings, current paradigms still express reasoning through textual continuation instead of visual or physical dynamics. In contrast, video represents reasoning as a process of visual continuation over time. Each frame in a video builds upon its previous ones, capturing the dynamics of motion, spatial consistency, and temporal causality within 2D and 3D space. The continuous and structured nature of frames makes video an ideal substrate for multimodal reasoning. Building on this insight, we propose reasoning via video, where reasoning emerges through next-frame generation rather than next-token prediction.
However, a comprehensive testbed for reasoning via video is lacking. To this end, we introduce VR-Bench, a dedicated benchmark designed to systematically assess the reasoning capabilities of video generation models. As shown in Figure 1, we ground our benchmark in the maze-solving task, a natural fit for visual reasoning due to its open-ended solution space and rich trajectory-based supervision. Each instance inherently demands spatial planning, dynamic tracking, and multi-step reasoning, making it an ideal testbed for evaluating model inference quality over time. Our dataset comprises 7,920 procedurally generated maze-centric videos, each paired with a corresponding Trace Reasoning Task that requires models to infer the optimal path. To ensure broad generalizability and challenge model robustness, VR-Bench spans five distinct maze types—Regular Maze, Irregular Maze, 3D Maze, Sokoban, and Trapfield—covering a wide spectrum of spatial structures and decision patterns. Additionally, each maze is rendered in diverse visual styles across more than a dozen themes, enabling fine-grained analysis of how well models generalize across varied visual domains and increasing the realism and complexity of the reasoning tasks.
Building upon the proposed VR-Bench, we conduct a systematic study of the reasoning via video paradigm. We construct instruction-following datasets derived from VR-Bench to elicit the reasoning capability of open-source video models. After supervised fine-tuning (SFT), these models exhibit significant performance gain across all reasoning tasks in VR-Bench. Moreover, SFT endows video models with strong out-of-domain generalization under diverse distribution shifts, including task difficulty, background style, and task type. Compared with vision–language models (VLMs) [comanici2025gemini, Qwen2.5-VL, li2024llava, jiang2025mme, chen2025mint] that reason via text, video models consistently outperform their counterparts on high-complexity reasoning tasks, showing greater stability and even superior performance as task difficulty increases, across diverse scenarios and tasks. This finding confirms that videos serve as a more expressive substrate for spatial reasoning, which facilitates video models to leverage temporal continuity and dynamic visual context. Interestingly, we further observe that video models exhibit a test-time scaling effect analogous to that of LLMs. As the inference budget increases, their performance improves substantially. By employing diverse sampling strategies at test time, video models effectively explore multiple reasoning trajectories, reducing uncertainty and achieving an average performance gain of 10–20%. These empirical results highlight the unique potential and scalability of the reasoning via video paradigm.
Our contributions are summarized as follows:
  • We make an early and systematic exploration of the reasoning via video paradigm, where reasoning emerges from sequential frame generation rather than token prediction. Compared with text-based approaches, this paradigm naturally captures temporal continuity and spatial causality, offering a more expressive and scalable substrate for solving spatial reasoning tasks.
  • We construct VR-Bench, a comprehensive benchmark grounded in maze-solving tasks with diverse spatial structures, difficulty levels, and texture styles. It provides fine-grained trajectory-level supervision and supports evaluations on path accuracy, rule compliance, generalization.
  • Through extensive experiments, we demonstrate that video-based reasoning outperforms text-based reasoning (e.g., VLMs) on complex tasks, especially under distribution shifts in maze type, visual style, and difficulty. Fine-tuned video models exhibit stronger performance, lower path redundancy, and higher structural fidelity.
  • We reveal a test-time scaling effect for video models, where performance consistently improves with larger inference budgets. Similar to that in LLMs, diverse sampling unlocks multi-path exploration and yields up to 20% performance gains across metrics and difficulty levels.
2 Related Works
2.1 Video Generation
Video models have advanced rapidly in both understanding and generation. Early understanding methods, such as MViT [fan2021multiscale], Video Swin Transformer [liu2022video], and VideoMAE [tong2022videomae], focused on learning robust video representations for downstream tasks. With LLMs [blattmann2023stable, ouyang2022training, anil2023palm], recent approaches tokenize videos and leverage language backbones for captioning [tong2025g], event localization [tian2018audio], and reasoning [hu2025video]. On the generation side, Sora-2 [brooks2024video] achieved controllable, physically consistent outputs with synchronized dialogue and sound. Proprietary systems such as Runway’s Gen-3 [runway2024gen3], Pika Labs [pikalabs2024pika15], Luma AI [lumalabs2024dreammachine], and Google DeepMind’s Veo series [deepmind2024veo2, deepmind2025veo3] further enhance video quality and realism but remain closed-source. In contrast, open-source frameworks such as Stable Video Diffusion [blattmann2023stable], OpenSora[opensora], Hunyan-Video [kong2024hunyuanvideo], and the Wan series [wan2025wan] democratize access, offering efficient architectures and scalable training for state-of-the-art video synthesis.
2.2 Evolution of Reasoning Paradigms
Chain-of-Thought (CoT) prompting has significantly enhanced the reasoning abilities of language models [wei2022chain, wang2022self, guo2025deepseek]. Reinforcement learning further integrates CoT-style reasoning into model training, enabling models to internalize multi-step thought processes. More recently, such paradigms have been extended to vision-language models (VLMs). Systems like o3 and o4-mini [openai2025o3systemcard] introduce the ”Think with Image” framework, where reasoning is grounded in visual operations such as zooming and cropping. This allows the model to dynamically interact with image regions as part of the CoT process, thereby improving multimodal reasoning [zheng2025deepeyes, su2025pixel, zhu2025active]. In parallel, the rise of unified models for both generation and understanding has given birth to a new reasoning paradigm centered on interleaved vision-language outputs. Instead of purely textual reasoning traces, these models generate coherent sequences that alternate between textual and visual elements [xie2025show, wu2025qwen, duan2025got, wan2025deepresearch], providing a more grounded and expressive format for complex multimodal reasoning.
2.3 Evaluation of Video Generation Reasoning
Previous benchmarks for video generation models have predominantly focused on assessing visual quality, temporal coherence, and alignment with human preferences [huang2024vbench, huang2024vbench++, han2025video, yuan2024chronomagic]. However, these evaluations largely neglect the reasoning capabilities of video models. Recent works have begun to explore reasoning via video generation—the ability of models to solve reasoning tasks through the generation process itself [guo2025video, wiedemer2025video, tong2025thinking]. For instance, models like Veo 3 demonstrate zero-shot competence in tasks such as maze navigation and symmetry recognition. These tasks require perceiving, modeling, and manipulating the visual world, indicating that video generation can inherently support spatial-temporal reasoning. Despite these promising directions, current benchmarks for video reasoning still suffer from several limitations: (1) Lack of fine-grained and objective evaluation: Current evaluations rely heavily on manual inspection or coarse metrics, without capturing the reasoning trajectory embedded in the video; (2) Absence of modality comparisons: There is a lack of systematic comparison with think with text or think with image paradigms, making it unclear whether video generation truly provides unique advantages for reasoning; (3) Neglect of tuning and scaling analysis: Unlike language or multimodal models, video reasoning benchmarks seldom explore whether supervised fine-tuning (SFT) or test-time scaling can improve performance. These gaps call for a new benchmark that evaluates not only generation quality but also the reasoning process in videos, using rigorous metrics, multimodal comparisons, and extensible settings.
3 VR-Bench
3.1 Dataset Construction
The VR-Bench dataset is a Visual Trace Reasoning (VTR) dataset that constructs various Maze Puzzles into visual reasoning tasks. Its construction process comprises two steps: Maze Generation and Video Generation.
Figure 2: Variations of difficulty level and maze texture
Maze Generation.
The maze generation process in our work encompasses five distinct types, with all 7,920 maze instances in the dataset generated programmatically through custom code [tong2025code2logic]. Each type is tailored to assess specific visual reasoning capabilities, as elaborated below:
Method EM ($\uparrow$) SR ($\uparrow$) PR ($\uparrow$) SD ($\downarrow$)
Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko
VLM
Gemini-2.5-pro 2.8 36.1 13.9 2.8 25.0 4.2 37.5 13.9 2.8 31.9 9.5 47.9 57.6 19.4 33.8 25.0 1.9 0.0 0.0 1.1
Gpt-5 high 13.9 31.9 18.1 0.0 23.6 69.4 33.3 27.8 1.4 34.7 11.8 43.6 53.7 23.7 35.6 31.0 2.1 6.9 20.0 0.5
Qwen2.5-VL-7B 0.0 1.3 0.0 0.0 1.4 1.4 6.9 1.4 1.4 2.8 6.5 12.4 14.3 11.3 7.8 300.0 26.7 66.7 80.0 1150.0
Qwen2.5-VL-7B-SFT 12.5 29.2 22.2 31.9 29.8 52.8 34.7 52.8 36.1 37.5 32.5 45.1 71.6 59.3 43.0 52.6 2.3 11.3 4.5 3.0
$\boldsymbol{\Delta}\!\uparrow$ +12.5 +27.9 +22.2 +31.9 +28.4 +51.4 +27.8 +51.4 +34.7 +34.7 +26.0 +32.7 +57.3 +48.0 +35.2 -247.4 -24.4 -55.4 -75.5 -1147.0
General Video Model
Closed-Source
Veo-3.1-fast 0.0 0.0 0.0 0.0 2.8 40.3 36.1 38.9 48.6 43.1 20.2 24.8 28.2 13.4 21.7 195.3 111.5 80.7 33.5 112.3
Veo-3.1-pro 0.0 4.2 1.4 0.0 0.0 47.2 36.1 59.7 50.0 37.5 24.6 33.9 39.1 18.0 21.4 140.7 94.5 85.4 40.1 141.8
Sora-2 1.4 5.6 0.0 0.0 4.2 75.0 72.2 83.0 37.5 43.1 45.1 45.7 46.6 19.3 27.4 302.9 187.0 145.1 92.4 138.7
kling-v1 0.0 0.0 0.0 0.0 0.0 2.8 0.0 1.4 27.8 12.5 6.3 8.8 10.4 11.7 9.0 25.2 69.7 356.1
Seedance-1.0-pro 0.0 2.8 2.8 0.0 0.0 75.0 45.8 59.7 77.8 13.9 12.8 35.8 42.7 23.6 17.1 162.3 143.4 99.1 84.4 241.9
MiniMax-Hailuo-2.3 0.0 1.4 2.8 0.0 0.0 68.1 40.3 70.8 55.6 45.8 23.2 24.2 30.3 20.3 15.5 464.0 170.0 90.9 50.1 165.5
Open-Source
Wan2.5-i2v-preview 0.0 2.8 4.2 0.0 0.0 58.3 26.4 77.8 24.5 22.4 14.3 21.8 34.4 24.5 17.1 378.4 281.8 73.2 119.9 278.0
Wan2.2-TI2V-5B 0.0 0.0 0.0 0.0 0.0 6.9 12.5 0.0 31.9 11.1 6.6 9.1 7.1 12.8 9.2 388.7 66.1 5.4 176.6
Ours
Wan-R1 33.3 56.9 38.9 65.3 4.2 76.4 69.4 100.0 100.0 69.4 60.6 71.6 79.1 93.5 44.3 10.3 2.4 3.9 3.9 10.2
$\boldsymbol{\Delta}\!\uparrow$ +33.3 +56.9 +38.9 +65.3 +4.2 +69.5 +56.9 +100.0 +68.1 +58.3 +54.0 +62.5 +72.0 +80.7 +35.1 -12.8 -25.1 -7.8 -100.1
Table 1: The five tasks of VR-Bench correspond to Base (Regular Maze), Irrg (Irregular Maze), Trap (TrapField), 3D (3D Maze), and Soko (Sokoban). The best and second-best results in each column are bolded and underlined, respectively. “–” indicates that the model produced no successful cases for the corresponding task, making SD undefined. $\Diamond$ denotes the base model of Wan-R1, for comparisons.
  1. 1.
    Regular Maze. We generate mazes with a grid-based layout to focus on the model’s ability to perceive basic maze structures and its path-finding and problem-solving competence, serving as a fundamental testbed for maze reasoning tasks.
  2. 2.
    Trapfield. This type transforms the ”walls” of traditional mazes into grid-shaped trap regions, reversing the logic from ”finding feasible paths” to ”avoiding traps”. Beyond altering the problem-solving logic, the more flexible movement space also challenges the model’s ability to plan and find optimal paths.
  3. 3.
    Irregular Maze. Moving away from regular block-shaped paths, we adopt curve-based path designs. This design prevents the model from relying on coordinate-based position encoding, thereby rigorously evaluating its pure visual perception of maze layouts. It also explicitly decouples visual reasoning from text-based reasoning, focusing on the video model’s capability to reason via video itself.
  4. 4.
    Sokoban. We modify the underlying rules of traditional mazes by introducing the ”Sokoban” task mechanism. Models need to comprehend and apply Sokoban logic on top of path finding, increasing task complexity and emphasizing the model’s ability to internalize and apply logic.
  5. 5.
    3D Maze. By extending the maze to a 3D space, we employ a stereoscopic structural design to test the model’s spatial perception ability in 3D environments and its cross-dimensional path reasoning capability.
Maze Variations.
To evaluate the generalization ability on the VTR task and enhance robustness in adapting to diverse maze scenarios, we introduce variations across two key dimensions: (1) Difficulty Level: We define three difficulty grades (Easy, Medium, and Hard) by adjusting the maze size (e.g., expanding from 5×5 to 7×7), modifying the number of maze branches, and adding obstacles; (2) Maze Texture: We vary the textures of maze obstacles, paths, and other components using textures generated via procedural methods and generative models, as shown in Figure 2, which exposes the policies to a broad visual distribution and mitigates overfitting to clean, synthetic environments.
Video Generation.
To generate solution videos from maze images, we use a Breadth-First Search solver to compute the optimal path for each maze. These paths are rendered into videos at 24 fps and standardized to 192 frames (8 seconds) by adjusting playback speed, producing consistent image–video pairs for training and evaluation.
3.2 Metric Design
We selected two different evaluation paradigms to comprehensively assess our task, as detailed below.
Path Matching.
To objectively and comprehensively evaluate the VTR task, we perform target tracking across each frame of the model-generated videos to record the motion trajectory of the target. By comparing and analyzing these trajectories against the optimal path for each task [xu2025visual], we propose the following four evaluation metrics.
  1. 1.
    Exact Match (EM)
    Defined as $\mathrm{EM}_{i}=\prod_{j=1}^{n_{i}}\mathbb{I}(\hat{v}_{ij}=v_{ij})$. This metric measures whether the model successfully generates the complete and correct trajectory that aligns with the shortest optimal valid path. One step of deviation from the optimal solution is considered incorrect.
  2. 2.
    Success Rate (SR)
    Defined as $\mathrm{SR}_{i}=\mathbb{I}\!\left(\mathbf{p}^{(gen)}_{\text{end}}\in\mathcal{B}_{\text{goal}}\right)$. SR measures whether the generated trajectory successfully reaches the designated goal region. It reflects the model’s capability to complete the task by arriving at the target position, with a value of $1$ indicating successful goal attainment and 0 indicating failure to reach the goal.
  3. 3.
    Precision Rate (PR)
    Defined as $\mathrm{PR}_{i}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[\prod_{k=1}^{j}\mathbb{I}(\hat{v}_{ik}=v_{ik})\right]$. PR quantifies the proportion of consecutively correct steps along the optimal path. It offers a softer metric than EM, reflecting the model’s ability to make steady, meaningful progress toward the complete correct trajectory.
  4. 4.
    Step Deviation (SD)
    Defined as $\mathrm{SD}_{i}=\frac{L^{(gen)}_{i}}{L^{(gt)}_{i}}-1$. SD quantifies the relative path-length redundancy of the generated trajectory, representing how much longer the model’s path is compared to the optimal one. A smaller SD indicates higher efficiency and closer adherence to the optimal solution.
Figure 3: Bad case visualization and VLM-as-judge schematic
Rule Compliance.
Not all generated videos faithfully depict the ball following the maze path to complete its task. As shown in Figure 2, we observed numerous failure cases during testing, including the ball “breaking through maze walls,” “disappearing and reappearing,” and “inconsistent maze layouts across frames.” To consistently evaluate the model’s adherence to spatial and physical rules [huang2024vbench++], we designed a prompt-based assessment protocol that examines five key dimensions: (1) Motion Continuity of the main subject, (2) Temporal Consistency of the subject, (3) Trajectory Rationality of the main subject, (4) Structural Consistency of the maze, and (5) Interactional Rationality of subject–maze interactions. Each generated video is interpreted by a VLM, which identifies and scores potential violations of these rules. Specifically, each dimension is assigned a binary score—0 if the behavior is deemed unreasonable and 1 if it is reasonable. The scores from all five dimensions are then aggregated to compute a unified VLM-score, ranging from 0 to 5, which provides a quantitative measure of the overall rule compliance in the generated videos.
Since Structural Consistency offers only a binary judgment of whether the maze layout has changed, we introduce Maze Fidelity (MF) to quantitatively measure the degree of structural consistency across frames, which defined as $\mathrm{MF}=\frac{1}{M}\sum_{i=1}^{M}\left(1-\frac{|{p:|I_{0}(p)-I_{i}(p)|>\tau}|}{N_{i}}\right)$.
Here, $M$ is the number of sampled frames; $I_{0}$ and $I_{i}$ denote the background regions of the first and the $i$-th frames; $\tau$ is the pixel-difference threshold; and $N_{i}$ is the number of valid overlapping pixels. MF quantifies background stability across frames, with higher values indicating better preservation of the static maze layout.
4 Experiment
Method MF ($\uparrow$) VLM-Score ($\uparrow$)
Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko
General Video Model
Closed-Source
Veo-3.1-fast 43.2 86.3 22.5 69.1 63.5 0.8 2.5 1.1 1.7 2.0
Veo-3.1-pro 80.5 89.3 82.2 73.4 95.8 2.5 2.8 1.5 2.0 2.5
Sora-2 96.5 97.2 97.2 95.4 95.2 3.9 4.1 3.9 3.3 3.2
kling-v1 55.5 73.8 54.4 84.9 72.9 1.4 2.5 1.9 2.8 2.8
Seedance-1.0-pro 87.7 97.2 64.3 86.3 83.0 3.0 3.8 2.4 2.7 3.2
MiniMax-Hailuo-2.3 92.4 93.4 91.4 94.9 93.3 3.5 3.4 3.2 3.7 3.6
Open-Source
Wan2.5-i2v-preview 69.2 74.8 70.6 82.7 90.4 1.2 1.8 2.6 2.6 2.9
Wan2.2-TI2V-5B 85.5 97.4 83.7 94.7 93.5 2.8 3.1 1.5 3.3 3.1
Ours
Wan-R1 91.2 98.1 93.3 95.7 94.1 4.2 4.3 4.4 4.0 4.1
-Wan2.2-TI2V-5B +5.7 +0.7 +9.6 +1.0 +0.6 +0.3 +0.2 +0.5 +0.7 +0.5
Table 2: MF and VLM-Score denote Maze Fidelity and the rule-compliance score evaluated by a VLM. The best and second-best results in each column are bolded and underlined.
To comprehensively evaluate the reasoning capability of video models, we conduct experiments on our proposed VR-Bench. We evaluate both state-of-the-art proprietary and open-source video models on this benchmark. To highlight the advantages of video models over traditional multimodal approaches, we also include representative VLMs in our evaluation. In addition, we fine-tune the open-source video model Wan2.2-TI2V-5B using VR-Bench to investigate their generalization ability on reasoning tasks. This allows us to assess whether such reasoning capabilities can emerge via supervised fine-tuning, and whether they can generalize across different settings.
4.1 Training Configurations
To investigate how well open-source video models can acquire and generalize reasoning abilities through fine-tuning, we trained Wan-R1 based on the proposed dataset.
Specifically, we used the first scene from each of the five game types in our benchmark. For each game, we created two training settings: one using only easy samples, and another using a mixture of easy, medium, and hard samples. In each case, the data was split into 80% for training and 20% for validation.
All models were fine-tuned using the Accelerate framework on A100 GPUs. We adopted a LoRA-based training strategy on the Wan2.2-TI2V-5B architecture, with a learning rate of 1e-4, image resolution of $\times512$, and video length of 193 frames. We applied LoRA (rank 32) to key attention and feedforward modules (q, k, v, o, ffn.0, ffn.2) of the Dit backbone. Each model was trained for 5 epochs, with a dataset repetition factor of 100.
4.2 Implementation Details of Baselines
We compare the Wan-R1 against a wide range of baselines: 1) 6 Closed-source video models: Veo-3.1-fast, Veo-3.1-pro [deepmind2025veo3], Sora-2 [openai2025sora2], Kling-v1 [KlingAI2025], Seedance-1.0-pro [Seedance10Pro], MiniMax-Hailuo-2.3 [MiniMaxHailuo23]. 2) 2 Open-source video models: Wan2.5-i2v-preview [AlibabaWan25], Wan2.2-TI2V-5B [wan2025wan]. 3) 3 VLMs: Gemini-2.5-pro [comanici2025gemini], Gpt-5 high [GPT5ChatGPT], Qwen2.5-VL-7B [Qwen2.5-VL].
Video Models.
To ensure fair comparison across different video models, we standardized input preprocessing and output postprocessing. Since most maze images are square ($1:1$), we applied black-border padding to satisfy model-specific resolution or aspect ratio requirements, followed by center cropping to restore the maze region. Generated videos were temporally aligned according to each model’s fixed or adjustable duration settings.
Veo-3.1 and Veo-3.1-Pro generate fixed 8-second videos in a $9:16$ aspect ratio, with black-border padding and post-cropping applied. Doubao outputs 10-second, $1:1$ videos, padding and cropping inputs smaller than $300\times 300$ pixels. Kling produces 10-second videos without further adjustments. MiniMax also yields 10-second videos at its default 768p resolution, using the same padding–cropping scheme for small inputs. Sora-2 generates 10-second, $9:16$ videos, followed by black-border padding and cropping to maintain maze integrity.
VLM.
Given an initial observation image $I_{0}$, the VLM predicts an action sequence $a_{\text{pred}}=[a_{1},\dots,a_{T}]$, representing its intended movements in the environment. The actions are sequentially executed in the simulator to verify trajectory validity against the optimal reference $a_{\text{opt}}$.
For each task type, we define its corresponding action space. In TrapField, Sokoban, and Regular Maze, actions correspond to four-directional moves {up, down, left, right}, and the VLM outputs sequences such as [“up”, “right”, “down”]. In the Irregular Maze, actions are defined over irregular graph nodes (A, B, …), where the VLM outputs node transition sequences like [“A”, “C”, “E”]; the model is evaluated on whether these transitions correspond to valid path connections. For the 3D Maze, actions include six directional movements covering both horizontal and vertical axes, and the VLM outputs sequences such as [“forward_left”, “up”, “forward_right”].
5 Insights and Discussions
Task EM SR PR SD
Easy Med. Hard Easy Med. Hard Easy Med. Hard Easy Med. Hard
Base 0.0 0.0 0.0 8.3 8.3 4.2 13.2 3.8 2.7 154.8
0.0
(+0.0)
4.2
(+4.2)
0.0
(+0.0)
83.3
(+75.0)
41.7
(+33.4)
58.3
(+54.1)
40.1
(+26.9)
26.1
(+22.3)
6.0
(+3.3)
28.2
(-126.6)
10.1
Irrg 0.0 0.0 0.0 29.2 4.2 4.2 13.4 8.1 5.8 48.3 39.3
83.3
(+83.3)
66.7
(+66.7)
54.2
(+54.2)
95.8
(+66.6)
87.5
(+83.3)
62.5
(+58.3)
88.0
(+74.6)
74.8
(+66.7)
68.4
(+62.6)
3.5
(-44.8)
7.8
3.1
(-36.2)
Trap 0.0 0.0 0.0 0.0 0.0 0.0 6.0 6.4 8.9
62.5
(+62.5)
0.0
(+0.0)
12.5
(+12.5)
100.0
(+100.0)
95.8
(+95.8)
62.5
(+62.5)
86.7
(+80.7)
43.7
(+37.3)
56.7
(+47.8)
2.1
5.9
3.5
3D 0.0 0.0 0.0 54.2 16.7 25.0 7.6 9.8 15.2 53.0 87.9 33.6
41.7
(+41.7)
0.0
(+0.0)
4.2
(+4.2)
100.0
(+45.8)
79.2
(+62.5)
83.3
(+58.3)
78.7
(+71.1)
47.4
(+37.6)
58.7
(+43.5)
4.6
(-48.4)
10.2
(-77.7)
13.8
(-19.8)
Soko 0.0 0.0 0.0 20.8 8.3 4.2 14.6 8.3 5.6 354.2 61.4
4.2
(+4.2)
0.0
(+0.0)
0.0
(+0.0)
83.3
(+62.5)
45.8
(+37.5)
33.3
(+29.1)
62.2
(+47.6)
12.5
(+4.2)
10.4
(+4.8)
18.8
(-335.4)
84.5
(+23.1)
16.1
Table 3: Difficulty generalization of Wan-R1 on VR-Bench. Each task block compares the baseline (Wan2.2-TI2V-5B) and the fine-tuned model (trained only on Easy level) across difficulty levels (Easy, Medium, Hard) and four metrics (EM, SR, PR, SD). Green indicates improvements, red indicates degradation, gray denotes no change or undefined cases.
Task Model EM SR PR SD
Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko Base Irreg Trap 3D Soko
Wan2.2-TI2V-5B Baseline 0.0 0.0 0.0 0.0 0.0 6.9 12.5 0.0 31.9 11.1 6.6 9.1 7.1 12.8 9.2 388.7 66.1 5.4 176.6
Regular Maze Fine-tuned
33.3
(+33.3)
5.6
(+5.6)
1.4
(+1.4)
0.0
(+0.0)
0.0
(+0.0)
76.4
(+69.5)
8.3
(-4.2)
88.9
(+88.9)
69.4
(+37.5)
30.6
(+19.5)
60.6
(+54.0)
22.7
(+13.6)
25.2
(+18.1)
13.7
(+0.9)
19.0
(+9.8)
10.3
(-378.4)
51.7
(-14.4)
3.8
12.7
(+7.3)
49.3
(-127.3)
Irregular Maze Fine-tuned
0.0
(+0.0)
56.9
(+56.9)
0.0
(+0.0)
0.0
(+0.0)
0.0
(+0.0)
11.1
(+4.2)
69.4
(+56.9)
52.8
(+52.8)
79.2
(+47.3)
12.5
(+1.4)
16.6
(+10.0)
71.6
(+62.5)
18.1
(+11.0)
16.8
(+4.0)
15.5
(+6.3)
35.8
(-352.9)
2.4
(-63.7)
6.9
9.0
(+3.6)
40.7
(-135.9)
3D Maze Fine-tuned
0.0
(+0.0)
5.6
(+5.6)
0.0
(+0.0)
65.3
(+65.3)
0.0
(+0.0)
38.9
(+32.0)
31.9
(+19.4)
30.6
(+30.6)
100.0
(+68.1)
20.8
(+9.7)
6.8
(+0.2)
20.6
(+11.5)
6.7
(-0.4)
93.5
(+80.7)
15.0
(+5.8)
108.2
(-280.5)
10.9
(-55.2)
8.8
3.9
(-1.5)
80.6
(-96.0)
Sokoban Fine-tuned
0.0
(+0.0)
1.4
(+1.4)
1.4
(+1.4)
0.0
(+0.0)
4.2
(+4.2)
0.5
(-6.4)
22.2
(+9.7)
18.1
(+18.1)
22.2
(-9.7)
69.4
(+58.3)
15.7
(+9.1)
23.7
(+14.6)
36.2
(+29.1)
15.7
(+2.9)
44.3
(+35.1)
46.3
(-342.4)
34.4
(-31.7)
39.9
20.1
(+14.7)
10.2
(-166.4)
Trapfield Fine-tuned
0.0
(+0.0)
0.0
(+0.0)
38.9
(+38.9)
0.0
(+0.0)
0.0
(+0.0)
93.1
(+86.2)
40.3
(+27.8)
100.0
(+100.0)
79.2
(+47.3)
6.9
(-4.2)
10.9
(+4.3)
12.9
(+3.8)
79.1
(+72.0)
14.7
(+1.9)
10.0
(+0.8)
57.5
(-331.2)
16.8
(-49.3)
3.9
11.4
(+6.0)
57.8
(-118.8)
Table 4: Comparison between the baseline (Wan2.2-TI2V-5B) and task-specific fine-tuned models across five game types (Base, Irreg, Trap, 3D, Soko). Each cell reports absolute performance and relative change over the baseline on four metrics: EM, SR, PR, and SD. Green indicates improvement, red indicates degradation, and gray denotes no change or undefined difference.
In this section, we discuss the observations and insights we draw from our comprehensive evaluation experiments.
· Wan-R1 Outperforms Prior Models on VR-Bench. As shown in Table 1, our method consistently achieves top performance across nearly all tasks and evaluation metrics, demonstrating both high accuracy and rollout efficiency. Notably, Wan-R1 attains a perfect SR of 100.0 on the Trap and 3D maze tasks, highlighting its robust success capabilities even in complex environments. Compared to its base model Wan2.2-TI2V-5B, Wan-R1 achieves a remarkable EM improvement of +65.3 on 3D, and reduces SD by 100.1 on Soko. These gains underscore the effectiveness of our fine-tuning strategy in enhancing both correctness and trajectory quality across diverse reasoning settings.
· Success Alone Does Not Guarantee Efficient Reasoning. Some models manage to complete tasks, but their rollouts remain highly inefficient. For instance, Sora-2 and MiniMax-Hailuo-2.3 achieve strong SR of 75.0 and 68.1 on the Base task, yet their corresponding SD values reach 302.9 and 464.0, revealing substantial path redundancy. Even more striking, the open-source VLM Qwen2.5-VL-7B produces a very low SR of 1.4, but still yields a high SD of 300.0, indicating unstable or erratic generation. As shown in Table 1, Wan-R1 achieves a comparable or better SR while reducing SD to just 10.3, demonstrating its ability to generate correct and efficient trajectories consistently.
· Reasoning via Video Outperforms Reasoning via Text. Under the same training data and settings, we fine-tune both the vision-language model (Qwen2.5-VL-7B) and the video model (Wan2.2-TI2V-5B). As shown in Table 1, the video-based model (Wan-R1) yields significantly larger gains across all metrics and tasks, especially in challenging settings like Trap and 3D. In contrast, Qwen2.5-VL-7B-SFT shows only moderate improvements. This highlights the advantage of reasoning via video in learning temporal reasoning and efficient path planning over static VLMs.
· Rule Compliance and Structural Fidelity. As shown in Table 2, our model Wan-R1 consistently achieves the highest VLM-Score across all maze types, and performs competitively in MF, ranking top-2 in most categories: 1) it consistently attains the highest VLM-Score across all maze types ($\geq$ 4.0), indicating superior rule-following behavior in motion continuity, physical plausibility, and environmental interactions; 2) it ranks among the top performers in MF, especially excelling on structurally complex mazes like Irreg and 3D. 3) it shows consistent improvements over its base model Wan2.2-TI2V-5B, confirming the effectiveness of our training paradigm in enhancing both visual stability and behavioral correctness.
· Reasoning via Video Scales Better. 1) As illustrated in Figure 4, model performance consistently declines with increasing maze difficulty on the Trap and Irre Maze tasks. In Easy settings, VLMs often match or even surpass state-of-the-art video models. However, as the maze complexity escalates, VLMs experience a sharper performance drop compared to video models. On large-scale hard mazes, even top-tier VLMs such as Gemini-2.5-Pro and GPT-5 are outperformed by leading video models like Sora-2 and Seedance-1.0-pro. 2) We attribute this trend to fundamental differences in the reasoning paradigms of the two model families. VLMs rely on encoding static visual observations into textual tokens and performing reasoning within a language-dominant latent space. As maze size increases, the number of visual tokens grows substantially, leading to context-length saturation and degraded long-horizon reasoning. In contrast, video models reason via visual dynamics, constructing a temporally grounded CoF that maintains spatial continuity across time. This video-centric reasoning mechanism preserves efficiency as scene complexity increases, since the number of visual tokens per frame remains stable, with visual tokens carrying significantly higher information density than textual tokens, a finding validated by DeepSeek OCR [wei2025deepseekocr] through optical context compression. Remarkably, models such as Sora-2 even exhibit improved performance in the Irreg Maze under higher difficulty levels, particularly in the SR metric. 3) These observations suggest that reasoning via video constitutes a more native and scalable paradigm for visual reasoning, enabling temporal–spatial problem solving that remains robust under increasing environmental complexity.
Figure 4: Model performance (PR and SR) on Irregular Maze and Trapfield across difficulty levels. Each curve represents a baseline, while the dashed and dotted lines indicate VLM and Video Model averages.
· Test-Time Scaling for Video Models. 1) Test-time scaling (TTS), exemplified by self-consistency [wang2023selfconsistency], has shown strong effectiveness in text-based reasoning tasks. Its key intuition is that complex reasoning problems often admit multiple valid solution trajectories, and sampling diverse paths increases the likelihood of converging to a correct answer. Maze-solving naturally shares this property: the solution space is open-ended, and multiple routes can lead to the goal. Motivated by this, we apply TTS to video models by perturbing the sampling noise to generate diverse rollouts, and evaluate performance using Pass@K, which selects the best solution among $K$ independent attempts. 2) Figure 5 shows the scaling behavior of Wan-R1 on the Irregular Maze benchmark. As $K$ increases from 1 to 16, the model achieves steady gains across all difficulty levels, with improvements of roughly 10–20% depending on metric and difficulty. On Easy mazes, performance rises almost monotonically and nears saturation at higher $K$. Medium difficulty shows clear early gains, often improving by 5–10% from $K=1$ to $K=4$, followed by continued smaller increases. Even on Hard mazes, where Pass@1 scores are lower, TTS still yields consistent upward trends, recovering solutions that single attempts often miss. 3) Overall, these results demonstrate that TTS significantly enhances video-model reasoning in VTR tasks. By initializing generation from different noise conditions, the model explores multiple solution pathways within the maze’s open-ended search space. This multi-path exploration effectively unlocks additional reasoning capacity, enabling video models to produce trajectories that are more accurate, more reliable, and consistently closer to the desired solution than those obtained under standard single-sample inference.
· Difficulty Generalization The results in Table 3 demonstrate the strong difficulty generalization capability of Wan-R1 across all five maze tasks. Although the model is fine-tuned only on the Easy level, it consistently delivers substantial improvements over the Wan2.2-TI2V-5B baseline on Medium and Hard mazes as well. This indicates that Wan-R1 does not simply memorize small or low-complexity layouts; rather, it internalizes a more principled and transferable reasoning procedure. The gains observed across unseen difficulty tiers show that fine-tuning on small mazes induces broad generalization: the model acquires a maze-solving strategy that scales to larger, more intricate structures without additional supervision. Such behavior reflects a deeper structural understanding of the environment and verifies that Wan-R1’s improvements stem from reasoning-pattern internalization rather than task-specific overfitting.
· Maze Type Generalization. As shown in Table 4, fine-tuning on a single game (e.g., Regular Maze, Trapfield, or 3D Maze) not only improves in-domain performance but also yields substantial gains on unseen games across all metrics (EM, SR, PR, SD). This highlights the emergence of transferable video reasoning capabilities. Notably, models fine-tuned on 3D Maze exhibit strong generalization, this overall transfer pattern underscores that training on complex 3D structures fosters general reasoning skills applicable to other maze types.
· Texture Generalization. Although fine-tuned only on the Raw skin of each game type, the model shows consistent and often substantial gains on unseen textures (Skin2 and Skin3). As shown in Table 5, this pattern holds across all five task domains. For example, in the Base task, Skin3—never encountered during training—improves by +23.6 in EM and +41.6 in PR. In the more challenging 3D Maze, the generalization is even stronger, reaching +50.0 EM and +78.5 PR on Skin3. These results demonstrate the model’s strong texture-level generalization and indicate that the learned spatio-temporal reasoning transfers robustly to new visual styles.
Figure 5: Performance on Irregular Maze using Wan-R1 under test-time scaling. Results are shown across different sampling numbers ($K\in{1,4,8,12,16}$) and difficulty levels (Easy, Medium, Hard).
Task EM SR PR SD
Raw Skin2 Skin3 Raw Skin2 Skin3 Raw Skin2 Skin3 Raw Skin2 Skin3
Base 0.0 0.0 0.0 6.9 4.2 2.8 6.6 4.6 9.4 388.7 11.7 14.9
33.3 1.4 23.6 76.4 38.9 68.1 60.6 12.3 51.0 10.3 19.0 9.0
(+33.3)
(+1.4)
(+23.6)
(+69.5)
(+34.7)
(+65.3)
(+54.0)
(+7.7)
(+41.6)
(-378.4)
(+7.3)
(-5.9)
Irreg 0.0 0.0 0.0 12.5 4.2 2.1 9.1 12.0 47.8 66.1 39.5 42.0
56.9 22.2 15.3 69.4 15.3 23.6 71.6 36.5 26.2 2.4 5.1 7.7
(+56.9)
(+22.2)
(+15.3)
(+56.9)
(+11.1)
(+21.5)
(+62.5)
(+24.5)
(-21.6)
(-63.7)
(-34.4)
(-34.3)
3D 0.0 0.0 0.0 0.0 4.2 4.2 7.1 8.9 8.3 73.6 59.3
65.3 43.1 50 100.0 97.2 100 93.5 83.8 86.8 3.9 6.4 5.9
(+65.3)
(+43.1)
(+50.0)
(+100.0)
(+93.0)
(+95.8)
(+86.4)
(+74.9)
(+78.5)
(-67.2)
(-53.4)
Soko 0.0 0.0 0.0 31.9 2.8 7.1 12.8 9.3 8.1 5.4
4.2 0.0 1.4 69.4 34.7 58.3 44.3 21.0 14.1 10.2 58.6 82.6
(+4.2)
(+0.0)
(+1.4)
(+37.5)
(+31.9)
(+51.2)
(+31.5)
(+11.7)
(+6.0)
(+4.8)
Trap 0.0 0.0 0.0 11.1 0.0 0.0 9.2 9.0 7.5 176.6
38.9 9.7 0.0 100.0 38.9 1.4 79.1 29.0 9.9 3.9 8.7 18.5
(+38.9)
(+9.7)
(+0.0)
(+88.9)
(+38.9)
(+1.4)
(+69.9)
(+20.0)
(+2.4)
(-172.7)
Table 5: Texture generalization under different skin. For each task, the table reports the baseline performance, the results after fine-tuning on the Raw texture, and the relative change across three texture conditions (Raw, Skin2, Skin3) on EM, SR, PR, and SD. Green denotes improvement, red denotes degradation, and gray indicates no change or undefined differences.
6 Conclusion
In this work, we take a step forward in evaluating whether video models can reason via video generation. We propose VR-Bench, a comprehensive benchmark grounded in maze-solving tasks to assess the spatial reasoning ability of video models. Our experiments demonstrate that fine-tuned video models exhibit strong spatial reasoning and consistently outperform leading vision-language models. Moreover, our analysis reveals a test-time scaling effect akin to self-consistency in language models, underscoring the scalable potential of video-based reasoning.
Limitations and Future Work. While VR-Bench provides a focused and rigorous testbed for spatial reasoning, it currently emphasizes maze-centric tasks. Future iterations of VR-Bench will explore broader and more challenging reasoning scenarios. For instance, we plan to incorporate Olympiad-level problem-solving tasks, such as solving complex physics or mathematics competition problems via video-based visual reasoning. In addition, we aim to support embodied reasoning settings where models are required to predict or simulate coherent action sequences within interactive environments.